Isosceles triangle has a base length of 30 and height of 8. What is the length of the other two sides?
Found 2 solutions by rfer, edjones:
Answer by rfer(16322)

15^2+8^2=c^2
225+64
289
sqrt 289=17 each side

Answer by edjones(8007)

The line that bisects the base and is perpendicular to it and ends at the apex of the triangle is the height. It divides the triangle into two right triangles, each with a base of 15 and and altitude of 8.
15^2+8^2=c^2 Pythagoras
225+64=289
sqrt(289)=17 length of each other side.
.
Ed